OpenStudy (anonymous):

A new cruise shipline has just alunched 3 new ships; the pacific paradise, the arribean paradise, and the mediteranin paradise. The carribean paradise has 20 more delxe stateroomsthan the pacific paradise. the mediteranian paradise has 31 fewer deluxe staterooms than four times the number of deluxe staterooms on the pacific paradise. Find the number of deluxe stateroomsfor each of the ships if the total number of deluxe staterooms for the three ships is 739.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

p + c + m = 739 c = p + 20 m = 4p - 31 now we sub p + (p + 20) + (4p - 31) = 739 --- combine like terms 6p - 11 = 739 -- add 11 to both sides 6p = 739 + 11 6p = 750 -- divide both sides by 6 p = 750/6 p = 125 c = p + 20 c = 125 + 20 c = 145 m = 4p - 31 m = 4(125) - 31 m = 500 - 31 m = 469 check... p + c + m = 739 125 + 145 + 469 = 739 739 = 739 (correct) pacific paradise has 125 deluxe staterooms carribean paradise has 145 deluxe staterooms mediteranian paradise has 469 deluxe staterooms